A shopkeeper sold 16 articles for total of $400 and made a profit of $48.00. How much did each article cost him?
Answer by robertb(5830)

Let x = cost of each article.
We use the formula Profit = Revenue - Cost.
Profit = 48, as mentioned.
Revenue = 400.
If there were 16 articles, the total cost would be 16x.
Thus, after substitution, we get
48 = 400 - 16x
==>16x = 400 - 48 = 352
==>x = 352/16
x = 22
Therefore each article cost him $22.